1. Bruce, an appliance salesman, earns a commission of $50 for each washing machine he sells and $100 for each refrigerator. Last month he earned $500 in commissions. Find all possibilities for the number of each kind of appliance he could have sold. 2.
Answer by jrfrunner(365)

let W=number of washing machines sold
let R=number of refrigerators sold
let T=total of W and R sold
--
quantity eq 1) W+R=T or W=T-R
earnings eq 2) 50*W+100*R=500
---
Substitute W=T-R into equation 2
50*(T-R)+100*R=500
50*T-50*R+100*R=500
50*T+50*R=500
T+R=10
R=10-T and since W=T-R=T-(10-T)=2T-10
--
since we need R>=0 then R=10-T>=0 or T<=10
also need W>=0 then 2T-10>=0 or T>=5
therefore 5<=T<=10, R=10-T, W=2T-10
---
T=5 then R=5, W=0
T=6 then R=4, W=2
T=7 then R=3, W=4
T=8 then R=2, W=6
T=9 then R=1, W=8
T=10 then R=0, W=10